OpenStudy (swag):

what is the surface area of a cylinder w. radius 3 & height 6 @terenzreignz

terenzreignz (terenzreignz):

Surface area of a cylinder is given by \[\Large A = 2\pi r h + 2\pi r ^2\]

OpenStudy (swag):

what would the SA be please tell me fastttt

terenzreignz (terenzreignz):

People usually use calculators for this sort of thing... This BETTER NOT be a quiz though -.- \[\Large A = 2\pi \times 3\times 6 + 2\pi \times 9=54\pi\approx169.65\]

OpenStudy (swag):

its not and so its 169.65?

terenzreignz (terenzreignz):

Should be.

OpenStudy (swag):

Ha ok I can stop rushing now. Thank you very much @terenzreignz I was on a call with my teacher and I had to learn this likeeee nowww in 2 seconds. Now i can actually review it.

terenzreignz (terenzreignz):

<whistles>

OpenStudy (anonymous):

Closed or open cylinder?

terenzreignz (terenzreignz):

Now you bring that up? -.- Oh well, wasn't specified...

OpenStudy (anonymous):

I was going to ask it earlier but got occupied, didn't know there was a time limit :)